Enough of a monoprotic weak acid is dissolved in water to produce a 0.0133 M solution. The pH of the resulting solution is 2.38. Calculate the Ka for the acid.


1
Expert's answer
2022-03-10T10:44:21-0500

HX = H+ + X-

C0 0.0133 0 0

"\Delta"C -x x x

[C] (0.0133-x) x x

Ka = [H+] * [X-] / [HX] = x * x / (0.0133-x) = x2/(0.0133-x);

pH = 2.38, therefore [H+] = 10-2.38 = 0.0042 M;

For reaction of dissociation: [H+] = [X-] = 0.0042 M;

Ka = (0.0042)2 / (0.0133-0.0042) = 1.76 *10-5 / 0.0091 = 1.93 *10-3 = 0.00193.

Answer: 1.93 *10-3 or 0.00193.